Sarah bought 8 folders and 3 rulers. Leah bought 3 folders and 8 rulers. If Sarah's total purchase was $1.25 more than Leah's, how much was each folder?
Answer by richwmiller(17219)

8f+3r=s
3f+8r=L
s-1.25=L
s=L+1.25
Can you take it from here?